OpenStudy (anonymous):

Given 3x^2 + x =1, find the value of 6x^3 - x^2 -3x + 2010

OpenStudy (radar):

Solve the quadratic 3x^2+x-1=0 first I try factoring well that didn't work so use the complete the square dividing thru by 3 getting x^2 + x/3 - 1/3= 0 add (1/6)^2 x^2 + x/3 + 1/36 = 1/3 +1/36 (x +1/6)^2 = 13/36 now take sq rt both sides x +1/6 = sq rt (13/36) \[x + 1/6 =\sqrt{13}/6\] \[x=-(1/6)\pm1/6\sqrt{13}\]

OpenStudy (radar):

Can you solve for x, and then substitute those values in your 2nd equation?

OpenStudy (radar):

\[x=-(1/6)\pm(1/6)\sqrt{13}\] That may be a little clearer.

OpenStudy (anonymous):

I will try to do the rest. Thanks...The Substitution part.

OpenStudy (radar):

Meantime I continue using calculator to get answer

OpenStudy (radar):

x=0.434 and x= -0.1

OpenStudy (radar):

I come up with 2009 and 2009.997 I know there is still another value but don't know how to solve for it. I also think that those answers can not be trusted.